Shift Summary for Eve Shift, 11/20/99

Ricardo_Barbosa (rbarbosa@fnal.gov)
Sun, 21 Nov 1999 00:22:00 -0600

Shift Summary for Eve Shift, 11/20/99

On Shift: Ricardo Barbosa, BobT

Hours of beam available at all intensities: 7:10
Hours of beam available at nominal intensity: 7:10
Hours of E799 data (special runs not included): 7:10
Hours of special runs (laser scans, muon runs...): 0:00

Accounting of lost time: Change the power suplly, DC4 trip off...

========================================================================

Runs ended this shift: 15020 and 15021

15020 ended at 15:35 with 19.0 million events;
409k online 3pi0-Dalitz)

Tapes brought to FCC this shift: UPA163-165
UPB163-165
UPC163-165
UPS055

========================================================================

Accesses during shift: We needed change a power suplly with problems.

========================================================================

Overview of shift:

========================================================================

Description of Shift:

run spill time comment

------------------------------------------------------------------------
15020 15:30 Ending run 15020. Tape4 @ 75.5%, Tape2 @ 95%.

15021 16:16 Starting run

19 16:33 DC4 trips off. Stopping trigger

28 16:44 Start trigger

46-47 Live time is very poor.

48 17:11 back to normal (probably the beam
position was moved, but now
is OK!)

52 daq dead for 2.8 sec

63 17:30 livetime not good (0.48)

84 18:00 DC4 trips off again...

91 18:10 Fixed(???) DC4 problem...

99 18:19 DC4 trips off again...we change the intensity
request to
9.5E12 until beam structure improves.

101 18:22 Fixed(???) DC4 problem...

109 18:32 DC4 trips off again...

111 18:35 Fixed(???) DC4 problem. We raise current trip to
be 80% of 1mAmp, was 80% of 100 uAmps.

164 19:45 Trigger stopped. Reseting the pipeline...

20:00 Discovered pipeline crate #10 has a failed
power supply, we call MCR for an access.

20:30 We attempt to stop run 15021, which hangs in
stop sequence. Status of EOF marks on tapes
unknown, we dismount tapes and fresh start a
new
run (15022).

21:00 Accesss complete, bad pwer supply for pipeline crate
#10 replaced.

15022 21:15 Started new run!


1-114 21:30- Bad Spill structure on the several-second
level.
23:55 MCR is struggling with it, but with limited
success.
TeV experts are investigating. Live time is
poor
during this period due to DAQ_DEAD and poor
instantaneous livetime.

========================================================================
Items to Follow up on Next Shift

Look the beam structure. A strange increase of intensity tripped off
DC4...

-- 
+---------------------------------------+
|					|
|	Ricardo Franco Barbosa		|
|		  diz			|
|		Ate' Logo!		|
|					|
+---------------------------------------+